How to calculate the autonomous spending

C = 200 + 0.75Y

Where:

Y = income

C = consumption

 

I = 250 I = investment

G = 400 G = government spending

T = 0.25Y T = taxes

X = 150 X = exports

M = 180 M = autonomous imports

Yf = 2 500 Yf = full employment income level 


1
Expert's answer
2021-03-23T08:35:18-0400


Total autonomous spending=C + I + G + X - M

Total autonomous spending = 200 +(0.75)(0) + 250 + 400 + 150 - 180 = 820